OpenStudy (anonymous):

A particle is moving on a straight line; at time t, its position is given by the rule s(t) = t^3-75t? Find the acceleration of the particle in ms^2 when the velocity is 0 t = secs s(t) = metres

OpenStudy (anonymous):

v(t) = s'(t) a(t) = v'(t) = s''(t) helps ?

OpenStudy (anonymous):

yes i have tried this not sure that i got the right answer !

OpenStudy (anonymous):

ok what did you get? when v = 0 ? and what is the acceleration at this time?

OpenStudy (anonymous):

s'(t) = v(t) = 3t^2 -75 find when v = 0 : 0 = 3t^2 - 75 t^2 = 25 -> t =5 s''(t) = v'(t) = a(t) = 6t a(5) = 30

OpenStudy (anonymous):

cheers thats what i got but went a bit further and confused and didnt need to cheers !! i had square root of 900 = 30ms^2 !!!

OpenStudy (anonymous):

:)